The maximum possible decrease in energy during the grain-growth of Copper, where the grain boundary energy is 0.5Jm^-2 and the initial grain diameter is 0.3mm is?
1
Expert's answer
2020-06-15T10:35:53-0400

The maximum possible decrease in energy is


"E=E_{gb}\cdot\frac{\pi d^2}{4}=0.5\cdot\frac{\pi 0.0003^2}{4}=3.53\cdot10^{-8}\text{ J}."